calculate moles of 100g of the electrons

Grade:12th pass

1 Answers

Sunil Kumar FP
9 years ago
mass of each electron=9.1*10^-31 kg
Therefore 9.1*10^-31 kg corresponds to 1 electron
100 gm of electron will correspond to =(1/9.1*10^-31)*(0.1) electron
=10^30 *.109
in terms of mole it will be=10^29 /(6.022*10^23)=1.66*10^5 mole